OpenStudy (cherokeegutierrez):

WILL GIVE MEDAL! What is the value of 2x^-2 y^-2 for x=3 and y=-2

OpenStudy (dinamix):

\[2(3) ^{-2}= 2*\frac{ 1 }{ 3^2 } = \frac{ 2 }{ 9 }\]

OpenStudy (dinamix):

\[y ^{-2}=(-2)^{-2}= \frac{ 1 }{ (-2)^2}=\frac{ 1 }{ 4}\] @CherokeeGutierrez

OpenStudy (anonymous):

plug in what you have for x and y into the apropiate spots